A boy on a skateboard coasts along at 5 m/s. He has a ball that he can throw at a speed of 10 m/s. What is the ball's speed relative to the ground if he throws the ball (a) forward or (b) backward?


1
Expert's answer
2021-04-26T17:13:35-0400

a) If the boy throws the ball forward, ball's speed relative to the ground is sum of speed of boy relative to the ground plus speed of the ball relative to the boy: "v_f = 5 \frac{m}{s} + 10 \frac{m}{s} = 15 \frac{m}{s}".

b) If the boy throws the ball backward, the speed is "v_b = 5 \frac{m}{s} - 10 \frac{m}{s} = -5 \frac{m}{s}", i.e. "5 \frac{m}{s}" in the backwards direction.